OpenStudy (anonymous):

Find the range of the function. f(x) = (x - 2)^2 + 2

OpenStudy (anonymous):

All real numbers y ≥ 0 y > 2 y ≥ 2 this one ?

hero (hero):

1. Expand the expression on the right side until we get a quadratic expression: f(x) = (x - 2)^2 + 2 = (x - 2)(x - 2) + 2 = x(x - 2) -2(x - 2) + 2 = x^2 - 4x - 4x + 4 + 2 = x^2 - 8x + 6 We know the graph opens upward, so next find the vertex using x = -b/(2a): x = -(-8)/2(1) = 8/2 = 4 y = (4)^2 -8(4) + 6 = 16 - 32 + 6 = -(32 -16 - 6) = -(32 - 22) = -10 So the vertex of the parabola is (4, -10) This means that vertically, the parabola begins at y = -10, and extends upward. Therefore, the range of the function is (-10, ∞)

hero (hero):

@lornbeach